# A Distributed Approach to Autonomous Intersection Management via Multi-Agent Reinforcement Learning

### v1

@@ Abstract
journal: Agents in Traffic and Transportation (ATT 2024)
@@ 1 Introduction
Connected autonomous vehicles (CAVs) [1] represent a groundbreaking advancement in transportation: with the potential to revolutionize mobility, CAVs are poised to redefine how we commute, park, travel, transport and interact with our urban environments [2, 3]. Equipped with advanced sensors and artificial intelligence systems [4], CAVs can navigate roads with precision, mitigating the risk of accidents caused by human error [5, 6]. This enhanced safety feature not only saves lives but also instills confidence in passengers, making mobility more accessible and inclusive for individuals with disabilities or vulnerabilities [7, 8, 9]. On a societal level, the advantages of CAVs extend far beyond individual convenience. By optimizing traffic flow and minimizing congestion, CAVs have the potential to alleviate the strain on highways [10], reducing travel times and improving overall efficiency and stability [11]. Such a traffic management not only benefits commuters but also enhances the productivity of cities and supports economic growth [12, 13]. In addition, CAVs hold the promise of environmental sustainability. With the integration of electric and hybrid propulsion systems, CAVs can significantly reduce greenhouse gas emissions [14, 15] and combat climate change contributing to a cleaner, greener connected future [16].
@@
A vast literature exists on AIM. The research in this field spans multiple fronts, each leveraging distinct methodologies to address the challenges of optimizing traffic flow and ensuring safety in dynamic urban environments. By employing reinforcement learning (RL), AIM systems can effectively learn and adapt intersection control strategies in response to changing traffic conditions [19, 20]. These systems typically comprise priority assignment models, intersection control model learning, and safe brake control mechanisms. Experimental simulations demonstrate the superiority of RL-inspired AIM approaches over traditional methods, showcasing enhanced efficiency and safety. Graph neural networks (GNNs) have also garnered attention for their potential in AIM [21, 22]. Leveraging RL algorithms, GNNs optimize traffic flow at intersections by jointly planning for multiple vehicles. These models encode scene representations efficiently, providing individual outputs for all involved vehicles. Game theory serves then as a foundational framework for MARL approaches in AIM. Indeed, game-theoretic models facilitate safe and adaptive decision-making for CAVs at intersections [23, 24]. By considering the diverse behaviors of interacting vehicles, these algorithms ensure flexibility and adaptability, thus enhancing autonomous vehicle performances in challenging scenarios. Finally, recursive neural networks (RNNs) integrated in the MARL framework represent an interesting approach in AIM research, with the objective of learning complex traffic dynamics and optimizing vehicle speed control [25].
@@
Despite the theoretical advancements in AIM techniques, their implementation still faces important challenges. The main obstacle is represented by the need for a highly expensive centralised server which has to be positioned in the proximity of the intersection, in order to simultaneously control all the vehicles. Moreover, the vehicles should continuously send their local information to this centralised controller, which will then gather and elaborate the data coming from all the road users, before sending back to each vehicle a velocity or acceleration command. Given the complexity and high demands of this technological framework, the integration of AIM devices into existing transportation infrastructures still requires many years of extensive research and testing. In this direction, we devise an alternative distributed approach based on the 3D surround view technology for advanced assistance systems [26]. As demonstrated in the sequel, such a method allows the reconstruction of a 360 scene centered around each CAV, which is useful to recover the information required for each agent involved in the proposed MARL-based technique. This, in turn, grants to effectively carry out AIM in a decentralised fashion, exploiting sensors that are currently available on the market, without the need for the centralised infrastructure described in the previous lines.
More precisely, the contributions of this paper are multiple.
@@
    Last but not least, we guarantee full reproducibility [^2] The Python code of our work can be found at https://github.com/mcederle99/MAD4QN-PS. The authors want to stress the fact that reproducibility represents a crucial issue within this field of research. Unfortunately, in many interesting related works we did not find reproducible code [19, 20, 25]. of the code that is used for the generation of the virtual experiments shown in this manuscript.
@@ 2.2 Q-Learning and Deep Q-Networks
where $\alpha>0$ is called the step-size parameter. The policy derived from the state-action value function is usually the $\varepsilon$-greedy policy, suitable to balance the trade-off between exploration and exploitation. [30]:
@@ 3.1 System modelling and design
$r=\begin{cases}+x\hskip 24.18501pt&\text{if}\hskip 8.5359pt||x||>0,\\ -k&\text{if vehicle not moving,}\\ -10\cdot k&\text{if a collision occurs,}\\ +10\cdot k&\text{if scenario completed,}\end{cases}$ (9)
@@ 4 Experiments on virtual environments
To develop our code Python 3.8 was employed along with the version 1.4 of the Deep Learning library PyTorch [40]. Moreover, a NVIDIA TITAN Xp GPU was used to run our experiments.
@@ 4.1 Baselines
    Two symmetric (N/S & W/E) fixed-time traffic lights (FTTL) [22, 25], with different cycle lengths.
@@
The parameters of the four traffic lights configurations are reported in Table 3.
@@ 4.2 Results
To assess the quality of MAD4QN-PS we consider four metrics, namely the travel time, the waiting time, the average speed and the collision rate. We remark that we have accounted for vehicle-centered metrics because of the decentralised nature of our algorithm. However, it is evident that by optimizing the performance of each single road user we also implicitly improve the quality of the whole intersection.
The robustness of our method is ensured by performing training ten times across different seeds, and then considering all the different trained models while evaluating our strategy. In particular, each model has been tested by running a cycle of the evaluation phase presented in Section 3.3, and then the obtained results from the different models have been averaged out. Finally, to ensure a fair comparison and analysis of the results, the same evaluation setup has been adopted for all the baselines introduced above.
@@
Figure 2(b) instead displays the average speed of each vehicle. The results shown in this histogram are clearly related to those in Figure 2(a); indeed, also in this case we can see that our method outperforms traffic lights control schemes. This occurs since the vehicles almost never stop, thus keeping a smoother velocity profile throughout all the duration of the simulation.
@@ 5 Conclusions and future directions
In future works, we aim to explore different directions for advancements. In particular, one of the main objectives is to also consider human driven vehicles inside the environment. Given the decentralised nature of the proposed method, we expect to render our algorithm more robust without dramatically change it. Conversely, significant redesign would be necessary for a centralised AIM approach. Furthermore, we envisage to test more complicated scenarios, both in terms of dimension and layout, again to improve the robustness of our algorithm. Finally, we intend to implement our algorithm in a scaled real-world scenario with miniature vehicles [43], to practically demonstrate the applicability of our method.

### v2

@@ Abstract
journal: ATT’24: October 19, 2024, Santiago de Compostela, Spain
@@ 1 Introduction
Connected autonomous vehicles (CAVs) represent a groundbreaking advancement in transportation, poised to revolutionize mobility by redefining commuting, parking, travel, and urban interaction [1, 2]. Equipped with advanced sensors and AI systems, CAVs navigate roads with precision, reducing accidents caused by human error [3, 4]. This enhanced safety feature not only saves lives but also makes mobility more accessible and inclusive for individuals with disabilities or vulnerabilities [5, 6]. On a societal level, CAVs optimize traffic flow and minimize congestion, reducing travel times and improving overall efficiency and stability [7, 8]. Additionally, CAVs promise environmental sustainability by integrating electric and hybrid propulsion systems, significantly reducing greenhouse gas emissions [9, 10].
@@
A vast literature exists on AIM. The research in this field spans multiple fronts, each leveraging distinct methodologies to address the challenges of optimizing traffic flow and ensuring safety in dynamic urban environments. By employing reinforcement learning (RL), AIM systems can effectively learn and adapt intersection control strategies in response to changing traffic conditions [13, 14]. These systems typically comprise priority assignment models, intersection control model learning, and safe brake control mechanisms. Experimental simulations demonstrate the superiority of RL-inspired AIM approaches over traditional methods, showcasing enhanced efficiency and safety. Graph neural networks (GNNs) have also garnered attention for their potential in AIM [15, 16]. Leveraging RL algorithms, GNNs optimize traffic flow at intersections by jointly planning for multiple vehicles. These models encode scene representations efficiently, providing individual outputs for all involved vehicles. Game theory serves then as a foundational framework for MARL approaches in AIM. Indeed, game-theoretic models facilitate safe and adaptive decision-making for CAVs at intersections [17, 18]. By considering the diverse behaviors of interacting vehicles, these algorithms ensure flexibility and adaptability, thus enhancing autonomous vehicle performances in challenging scenarios. Finally, recursive neural networks (RNNs) integrated in the MARL framework represent an interesting approach in AIM research to learn complex traffic dynamics and optimize vehicle speed control [19].
@@
Despite the advancements in AIM techniques, their implementation still faces important challenges. One of the main obstacles is represented by the need for an expensive centralised server which has to be positioned in the proximity of the intersection, in order to simultaneously control all the vehicles. Moreover, the vehicles should continuously send their local information to this centralised controller, which will then gather and elaborate the data coming from all the road users, before sending back to each vehicle a velocity or acceleration command. Given the complexity and high demands of this technological framework, the integration of AIM devices into existing transportation infrastructures still requires many years of extensive research and testing. In this direction, we devise an alternative distributed approach based on the 3D surround view technology for advanced assistance systems [20]. As shown in the sequel, such a method allows the reconstruction of a 360 scene centered around each CAV, which is useful to recover the information required for each agent involved in the proposed MARL-based technique. This, in turn, grants to effectively carry out AIM in a decentralised fashion, exploiting sensors that are currently available on the market, without the need for the centralised infrastructure described in the previous lines. More precisely, the contributions of this paper are multiple.
@@
    Last but not least, we guarantee full reproducibility [^2] The Python code of our work can be found at https://github.com/mcederle99/MAD4QN-PS. The authors want to stress the fact that reproducibility represents a crucial issue within this field of research. of the code that is used for the generation of the virtual experiments shown in this manuscript.
@@ 2.2 Q-Learning and Deep Q-Networks
where $\alpha>0$ is called the step-size parameter. The policy derived from the state-action value function is usually the $\varepsilon$-greedy policy, suitable to balance the trade-off between exploration and exploitation [24]:
@@ 3.1 System modelling and design
$r=\begin{cases}+x\hskip 24.18501pt&\text{if}\hskip 8.5359pt{\color[rgb]{0,0,0}x}>0,\\ -k&\text{if vehicle not moving,}\\ -10\cdot k&\text{if a collision occurs,}\\ +10\cdot k&\text{if scenario completed,}\end{cases}$ (9)
@@ 4.1 Baselines
    Three symmetric (N/S & W/E) fixed-time traffic lights (FTTL), considering two cycle lengths already analyzed in [16, 19], and also the optimal cycle length in function of the traffic flow, computed according to Webster’s formula [37]. The final flow rate of vehicles for evaluation has been set to 600 veh/hour, as will be discussed in Section 4.2.
@@
The parameters of the five traffic lights configurations are reported in Table 3.
@@ 4.2 Results
[…]
@@ 5 Conclusions and future directions
In future works, we aim to explore different directions for advancements. In particular, one of the main objectives is to also consider human driven vehicles inside the environment and extend our approach to this field of research (see, e.g. the initial effort made in [38]). In this case, the most challenging issue is indeed represented by the synchronization of traffic lights accounting for the presence of human driven vehicles. Moreover, given the decentralised nature of the proposed method, we expect to render our algorithm more robust without dramatically change it. Conversely, significant redesign would be necessary for a centralised AIM approach. Furthermore, we envisage to test more complicated scenarios, both in terms of dimension and layout, again to improve the robustness of our algorithm. Finally, we intend to implement our algorithm in a scaled real-world scenario with miniature vehicles [39], to practically demonstrate the applicability of our method.